# Point cloud-based diffusion models for the Electron-Ion Collider

### v1

@@ III.1 Kinematic distributions with full PID
Next, we consider kinematic variables that are particularly relevant for the analysis of electron-proton scattering data. First, we consider the Deep Inelastic Scattering (DIS) cross-section, which is differential in the scaling variable Bjorken $x_{B}$ and the photon virtuality
$x_{B}=\frac{Q^{2}}{2P\cdot q}\,,\quad Q^{2}=-q^{2}=-\left(k-k^{\prime}\right)^{2}\,.$ (6)
@@
Figure 5: Top row: Diffusion model results and comparison to Pythia8 for the DIS variables Bjorken $x_{B}$ and photon virtuality $Q^{2}$. Bottom row: Analogous comparison for the energy $z_{h}$ of pions and transverse momentum $q_{T}$ in the Breit frame relevant for SIDIS.
@@ IV Conclusions and Outlook
The code for this paper can be found at https://github.com/ViniciusMikuni/OmniLearn.
@@ Acknowledgements.
We would like to thank Kaori Fuyuto, Chris Lee, and Emanuele Mereghetti for helpful discussions. JYA, FR, and NS were supported by the U.S. Department of Energy, Office of Science, Contract No. DE-AC05-06OR23177, under which Jefferson Science Associates, LLC operates Jefferson Lab. JYA and FR were supported in part by the DOE, Office of Science, Office of Nuclear Physics, Early Career Program under contract No DE-SC0024358. NS and RW are supported by the DOE, Office of Science, Office of Nuclear Physics in the Early Career Program. VM and FT are supported by the U.S. Department of Energy (DOE), Office of Science under contract DE-AC02-05CH11231. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility using NERSC award NERSC DDR-ERCAP0030239. This research was supported in part by the Quark-Gluon Tomography (QGT) Topical Collaboration, under contract no. DE-SC0023646.

### v2

@@ III.1 Kinematic distributions with full PID
Next, we consider kinematic variables that are particularly relevant for the analysis of electron-proton scattering data. First, we consider the Deep Inelastic Scattering (DIS) cross-section, which is differential in the scaling variable Bjorken $x$ and the photon virtuality
$x=\frac{Q^{2}}{2P\cdot q}\,,\quad Q^{2}=-q^{2}=-\left(k-k^{\prime}\right)^{2}\,.$ (6)
@@
Figure 5: Top row: Diffusion model results and comparison to Pythia8 for the DIS variables Bjorken $x$ and photon virtuality $Q^{2}$. Bottom row: Analogous comparison for the energy $z_{h}$ of pions and transverse momentum $q_{T}$ in the Breit frame relevant for SIDIS.
@@ IV Conclusions and Outlook
The code for this paper can be found at https://github.com/ViniciusMikuni/OmniLearn while the data generated to train the model is available at https://zenodo.org/records/14027110.
@@ Acknowledgements.
We would like to thank Kaori Fuyuto, Chris Lee, Emanuele Mereghetti and Benjamin Nachman for helpful discussions. JYA, FR, and NS were supported by the U.S. Department of Energy, Office of Science, Contract No. DE-AC05-06OR23177, under which Jefferson Science Associates, LLC operates Jefferson Lab. JYA and FR were supported in part by the DOE, Office of Science, Office of Nuclear Physics, Early Career Program under contract No DE-SC0024358. NS and RW are supported by the DOE, Office of Science, Office of Nuclear Physics in the Early Career Program. VM and FT are supported by the U.S. Department of Energy (DOE), Office of Science under contract DE-AC02-05CH11231. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility using NERSC award NERSC DDR-ERCAP0030239. This research was supported in part by the Quark-Gluon Tomography (QGT) Topical Collaboration, under contract no. DE-SC0023646.